# DIO: Refining Mutual Information and Causal Chain to Enhance Machine Abstract Reasoning Ability

### v1

@@ VIII DIEGO: Methods for Rectifying the Causal Chain in DIO
This study has made two attempts to refine the learning objectives of the DIO model, resulting in the development of two distinct methodology: Brando and WORLD. In this section, we devise a methodology named DIEGO (Direct Intervention for Enforcing Guided Optimal Reasoning), which is designed to rectifying the causal chains modeled by DIO, thereby advancing its reasoning performance.
@@ VIII-A Mutual Information Excels at Problem-Solving Rather Than Reasoning
This paper realizes that setting the maximization of the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$ as the learning objective of DIO is suboptimal. While mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$ directly correlates with accuracy in solving RPM problems, it fails to capture causal relationships inherent in the task design. This creates a critical discrepancy: DIO is optimized for high-accuracy RPM solution generation rather than learning the human-predefined reasoning logic. Both the Brando and WORLD methodologies aim to assist DIO in better maximizing mutual information, yet they fail to address this discrepancy.
Therefore, this paper argues that the causal chain underlying RPM problems should be further modeled. The network architecture of DIO strongly drew reference from the causal chain within RPM problems, with the hope that DIO can autonomously model the causal chain during the training process. However, this can only ensure that DIO is structurally aligned with the causal chain in RPM problems. Whether its latent variables are semantically aligned cannot be guaranteed and remains to be verified.
In order to verify semantic alignment, this paper selects the “Distribute 9” database under the RAVEN problems[19] as the research object. We conduct training on this database using DIO, as well as DIO supplemented with Brando and DIO supplemented with WORLD, respectively. Subsequently, we select 30 instances with strictly identical progressive patterns from this database and verify the pairwise similarity of the progressive pattern representations encoded by DIO for these 30 instances. The calculated $30\times 30$ similarity matrices have been plotted into heatmaps as shown in Figure 9.
@@
If DIO accurately models the causal chain embedded in RAVEN problems, then after these 30 instances with the same progressive patterns are encoded into progressive pattern representations by DIO, they should demonstrate high cosine similarity among themselves. However, the three similarity matrices in Figure 9 do not corroborate this expectation. This indicates that, even with auxiliary support from the Brando or WORLD methods, the causal chain modeled by DIO remains systematically misaligned with the ground-truth chains inherent in RPM problems. In addition, this further points out the limitations of using mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$ maximization as the learning objective. Given this, this paper proposes that rectifying the causal chain modeled by DIO during its training process will substantially improve performance in solving RPM problems.
@@
In preliminary experiments, we found that DIO exhibits deviations in modeling the segment of the causal chain that involves “abstract attributes $\rightarrow$ progressive patterns of attributes”. This corresponds to a misalign between the results $P_{\alpha{jkl}}$ derived from Formula (IV-C) and human semantics. Therefore, this paper proposes that refining the progressive pattern representations $P_{\alpha{jkl}}$ encoded by DIO would be an effective solution.
This paper has taken note of the metadata provided for each instance in RPM problems, which describes the corresponding instance’s progressive pattern. We argue that incorporating this metadata as a supervisory signal into DIO’s training process to supervise the derivation of $P_{\alpha{jkl}}$ can achieve the objective of rectifying the causal chain modeled by DIO. Thus, we have designed a method named DIEGO. DIEGO involves introducing an additional set of trainable reference vectors $\{V_{f}|f\in[1,F]\}$ for DIO, along with an auxiliary loss function term $\ell_{\text{DIEGO}}$.
@@
Specifically, the DIEGO method first establishes a set of trainable reference vectors $\{V_{f}|f\in[1,F]\}$, where each component represents a distinct metadata modality. Consequently, the reference vector set $\{V_{f}|f\in[1,F]\}$ can serve as a comprehensive dictionary for metadata. The loss term $\ell_{\text{DIEGO}}$ serves to align each progressive pattern representation $P_{\alpha{jkl}}$ extracted by DIO with its corresponding reference vector within $\{V_{f}|f\in[1,F]\}$ according to the accompanying metadata. The loss function $\ell_{\text{DIEGO}}$ can be expressed as follows:
@@
${\ell_{\text{DIEGO}}}=-\sum_{\hat{f}=1}^{F}\sum_{j=1}^{N}\sum_{k=1}^{9}\sum_{l=1}^{2}\text{Meta}_{{\hat{f}l}}\cdot\log\frac{{{e^{({P_{\alpha jkl}}\cdot{V_{\hat{f}}})/|\tau|}}}}{\sum_{f=1}^{F}{{e^{({P_{\alpha jkl}}\cdot{V_{f}})/|\tau|}}}}$ (43)
Here, $\text{Meta}_{\hat{f}l}$ denotes the indication status of the $l$-th metadata entry with respect to the $\hat{f}-$th type of progressive pattern. Meanwhile, $\tau$ represents a trainable positive scaling coefficient. The detailed expression of the loss term $\ell_{\text{DIEGO}}$ demonstrates our methodology for achieving alignment. This method utilizes a cross-entropy loss function with a scaling coefficient to control the cosine similarities between $P_{\alpha_{jkl}}$ and all components in the metadata dictionary $\{V_{f}|f\in[1,F]\}$ according to the indication information from the metadata.
When training the DIO model using $\ell_{\text{DIO}}+\ell_{\text{DIEGO}}$ as the loss function, a solving model that accurately captures the causal chains in RPM problem will be obtained. Previously, some studies have also attempted to treat metadata as additional supervision signals and introduce extra progressive pattern learning tasks into their solving models, aiming to simultaneously enhance the model’s interpretability and reasoning accuracy. However, research led by MRNet[25, 28, 27, 31] indicates that this approach may backfire. Specifically, if metadata is directly incorporated into the training process of the solving models they designed, the model’s reasoning performance will instead decline. This may be logically flawed, but it is indeed an established fact. Nevertheless, this paper argues that the DIO model constructed based on causal chains, coupled with the DIEGO method capable of strategically correcting key segments of the constructed causal chains, holds the key to breaking through this impasse.
@@ IX Solving Generative RPM Problems
Generative RPM problems represent an advanced variant of RPM tasks, where the solving model is required to generate an ideal answer for the context matrix $\{x_{i}|i\in[1,8]\}$ instead of selecting from a predefined bounded option pool.
IX-A DIO And WORLD Solve Generative RPM Problems
The WORLD method not only enhances the performance of DIO but also equips it with the capability to solve generative RPM problems. Specifically, the WORLD method introduces an external GMM $\{U_{m}|m\in[1,M]\}$ and a decoder $\mathcal{D}(\cdot)$ for DIO. Meanwhile, under the WORLD framework, the loss function term $\ell_{\text{Weighted}}$ defined in Formula (42) not only tightens the variational lower bound on the mutual information $I(\{x_{i}\}_{i=1}^{8};x_{\alpha})$ for DIO but also endows DIO with the ability to assign weights to each component in the GMM $\{U_{m}|m\in[1,M]\}$ with respect to the given context matrix features $\{z_{ij}|i\in[1,8],j\in[1,N]\}$. The weights assigned to the component $U_{m}$ for the contextual features $\{z_{ij}|i\in[1,8]\}$ can be calculated as follows:
$S_{{m}{j}}=\tilde{\text{DIO}}(\{z_{i{j}},U_{{m}}|i\in[1,8]\})$ (44)
Based on all computed weights $\{S_{mj}|m\in[1,M],j\in[1,N]\}$, we employ the Gumbel-Softmax trick to draw $N$ components from the GMM $\{U_{m}|m\in[1,M]\}$. This process can be formulated as:
@@ IX-A DIO And WORLD Solve Generative RPM Problems
$\hat{U}_{j}$ $=U_{\hat{m}_{j}}$ (48)
These $N$ components $\{\hat{U}_{j}|j\in[1,N]\}$ can be regarded as the distribution of solutions estimated by DIO for the context matrix $\{x_{i}|i\in[1,8]\}$. The decoder $\mathcal{D}$ is capable of transforming the samples from these $N$ components into the generated answer $\hat{x}$:
@@
Formulas (34), as well as formulas (44) - (49), indicate that the DIO trained by the WORLD method has the capability to estimate the distribution of solution for the context matrix $\{x_{i}\mid i\in[1,8]\}$ and generate RPM images that can serve as solutions.
IX-B DIEGO Validates the Rationality of Mutual Information Supervision
Verifying the solution accuracy of generative RPM problems has long been challenging. RPM problems often have multiple valid solutions, which means that the generative solving models cannot be evaluated simply by comparing each generated solution with the given correct option. This also makes it challenging for RPM experts to devise a perfect verification algorithm. Most prior work on generative RPM problems has used self-verification, i.e., using their own discriminative models to validate their own generative models.
However, this paper argues that this paradigm of self-verification is inadequate. The cause of the multiple-solution characteristic in RPM problems lies in the fact that solution identification is driven by a progressive pattern. Any RPM image that adheres to the progressive pattern exhibited in the contextual matrices can be considered a valid solution. This implies that algorithms or models used to verify the solution accuracy of generative RPM problems should also be progressive-pattern-driven. The DIEGO approach meets this expectation, as the DIO model trained solely with the loss term $\ell_{\text{DIEGO}}$ will be a progressive-pattern-driven solving model.
Additionally, DIO+WORLD is a strong mutual-information-based inductive model. Using the progressive-pattern-driven model DIEGO to verify DIO+WORLD’s ability to solve generative RPM problems not only avoids the imprecision of symmetric verification but also tests the deeper potential of mutual-information-driven solution models.
@@ X Experiment
All experiments conducted in this paper were programmed in Python, utilizing the PyTorch[37] framework. This paper introduces a comprehensive framework for solving RPM problems through three distinct configurations that integrate a purpose-built deep learning model with specialized training methodologies. The proposed configurations leverage:
  1. 1.
    Solo DIO, which solves RPM problems through a structured causal chain modeling approach without additional optimization strategies. In the DIO model, the hyperparameter $N$ is set to 16, indicating that the ViT within its image feature extraction module employs a $4\times 4$ patch division strategy;
  2. 2.
    Brando-enhanced DIO, which employs trainable hypothetical incorrect options to tighten the variational lower bound of mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$. In the Brando framework, the hyperparameter $\beta$ is set to 24, which means the Brando network provides 88 ($\beta-17=8$) constructive hypothetical incorrect options to the DIO in each iteration;
  3. 3.
    WORLD-augmented DIO, utilizing a Gaussian Mixture Model for feature distribution modeling to provide an infinite set of weighted incorrect options for tighting the lower bound of $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$. In the WORLD framework, the hyperparameter $M$ is set to 256, equipping the GMM employed for DIO with 256 components. This configuration enables the generation of $M^{N}=256^{16}$ distinct solution distributions, theoretically allowing for the sampling of an infinite diversity of incorrect options.
  4. 4.
    DIEGO-guided DIO, incorporating metadata supervision to directly rectify semantic biases in the causal chain of abstract attribute patterns. In the DIEGO framework, the parameter $F$ is determined by the number of metadata modality in the targeted RPM problem, rather than being arbitrarily set.
@@ X-A Experiment on RAVEN and I-RAVEN Problems
Experimental results recorded in Table I indicate that all four model configurations proposed in this study outperform all existing baselines on both the RAVEN and I-RAVEN datasets. Notably, the DIO+DIEGO configuration achieves a peak accuracy of 99.9%, delivering significant performance gains over prior state-of-the-art models (Triple-CFN[33] and RS-TRAN[31]). Particularly in complex reasoning tasks such as $\times3$ Grid and O-IG, the DIO series models demonstrate distinct advantages, validating the generalizability and robustness of their architectural design for abstract visual reasoning challenges.

### v3

@@ VIII DIEGO: Methods for Rectifying the Causal Chain in DIO
This study has made two attempts to refine the learning objectives of the DIO model, resulting in the development of two distinct methodologies: Brando and WORLD. In this section, we devise a methodology named DIEGO (Direct Intervention for Enforcing Guided Optimal Reasoning), which is designed to rectify the causal chains modeled by DIO, thereby advancing its reasoning performance.
@@ VIII-A Mutual Information Excels at Problem-Solving Rather Than Reasoning
This paper realizes that setting maximization of the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$ as the learning objective of DIO is suboptimal. Although this mutual information measure strongly correlates with the solving accuracy of RPM tasks, it neglects the human-predefined reasoning logic, yielding a model optimized for high accuracy rather than for faithful causal reasoning. Both the Brando and WORLD methodologies aim to assist DIO in better maximizing mutual information, yet neither resolves this limitation. In a nutshell, mutual-information-driven approaches remain effective but must be complemented by an explicit causal perspective.
Therefore, this paper argues that the causal chain underlying RPM problems should be further modeled. The network architecture of DIO is designed to mirror this causal chain, aiming for DIO to automatically capture the chain during training. Nevertheless, while this design guarantees structural alignment, it cannot ensure that the latent variables derived by DIO are also semantically aligned. Given that mutual information fails to account for causality, the semantic alignment awaits empirical verification.
In order to verify semantic alignment, this paper selects the “Distribute 9” dataset from the RPM-style RAVEN benchmark [19] as the research object. We conduct training on this dataset using DIO, as well as DIO supplemented with Brando and WORLD. Subsequently, we select 30 instances with strictly identical progressive patterns from this dataset and verify the pairwise similarity of the progressive pattern representations encoded by DIO for these 30 instances. The calculated $30\times 30$ similarity matrices are visualized as heatmaps in Figure 9.
@@
If DIO accurately models the causal chain embedded in RAVEN problems, then once DIO encodes 30 instances sharing the same progressive patterns into pattern representations, those representations should exhibit high cosine similarity among themselves. Contrary to expectations, the three similarity matrices in Figure 9 fail to support this anticipated outcome, indicating that, even with auxiliary support from the Brando or WORLD methods, the causal chain modeled by DIO systematically deviates from the ground-truth causal chains inherent in RPM problems. This underscores the limitations of adopting mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$ as the learning objective. Given these findings, we propose that further modeling of the causal chain could start with semantic rectification in DIO, which would substantially enhance its performance in solving RPM problems.
@@
In preliminary experiments, we found that DIO exhibits deviations in modeling the segment of the causal chain that involves “abstract attributes $\rightarrow$ progressive patterns of attributes”. This corresponds to a misalignment between the outputs $P_{\alpha{jkl}}$ produced by Formula (IV-C) and human semantics. Therefore, this paper proposes rectifying the progressive pattern representations $P_{\alpha{jkl}}$ encoded by DIO as an effective remedy.
This paper notes that each RPM instance is accompanied by metadata describing its progressive patterns that precisely match the semantics predefined for $P_{\alpha{jkl}}$ in the DIO architecture. We argue that incorporating this metadata as a supervisory signal into DIO’s training process to guide the derivation of $P_{\alpha{jkl}}$ can rectify the causal chain modeled by DIO.
@@
To this end, we propose the DIEGO method. DIEGO constructs a set of trainable reference vectors $\{V_{f}|f\in[1,F]\}$. By assigning each vector a distinct metadata modality, this entire set can be considered as a comprehensive metadata dictionary. In addition, DIEGO introduces an auxiliary loss term $\ell_{\text{DIEGO}}$ that serves to align each progressive pattern representation $P_{\alpha{jkl}}$ extracted by the DIO with its corresponding reference vector in $\{V_{f}|f\in[1,F]\}$ based on the accompanying metadata. This term is formulated as follows:
@@
${\ell_{\text{DIEGO}}}=-\sum_{\hat{f}=1}^{F}\sum_{j=1}^{N}\sum_{k=1}^{9}\sum_{l=1}^{2}\text{Meta}_{{\hat{f}l}}\cdot\log\frac{{{e^{\frac{\left({P_{\alpha jkl}}\cdot{V_{\hat{f}}}\right)}{|\tau|}}}}}{\sum_{f=1}^{F}{{e^{\frac{\left({P_{\alpha jkl}}\cdot{V_{f}}\right)}{|\tau|}}}}}$ (43)
Here, $\text{Meta}_{\hat{f}l}$ is a binary indicator that flags whether the $l$-th metadata entry matches the modality associated with the $\hat{f}-$th vector within the reference dictionary $\{V_{f}|f\in[1,F]\}$. Meanwhile $\tau$ is a trainable scaling coefficient. The detailed formulation of $\ell_{\text{DIEGO}}$ elucidates our alignment mechanism. Guided by the metadata-derived indicators, this loss term employs logarithmic smoothing and exponential scaling to regulate the cosine similarities between each $P_{\alpha_{jkl}}$ and all vectors in the reference dictionary. Evidently, our objective has shifted from statistical quantity control to vector similarity control, distinguishing DIEGO from Brando and WORLD: it is no longer mutual-information-driven but progressive-pattern-driven. When the DIO model is trained with the composite loss $\ell_{\text{DIO}}+\ell_{\text{DIEGO}}$ and the trainable reference dictionary $\{V_{f}|f\in[1,F]\}$, the forward pass of the relevant modules proceeds as illustrated in Figure 10.
Figure 10: Feedforward Process of DIO Model Incorporating $\ell_{\text{DIEGO}}$ as an Additional Loss Term.
After training, a solving model that captures causal chains in RPM problems is obtained.
It is widely accepted that introducing more supervisory signals should improve model performance. Indeed, previous studies have attempted to leverage metadata as additional supervision by designing progressive pattern-learning tasks, aiming to boost both interpretability and inference accuracy simultaneously. Recent work [25, 28, 27, 31] has collectively shown, however, that this practice can backfire: directly embedding metadata into proposed models actually degrades reasoning performance. This paper argues that the DIO model, constructed upon causal chains, together with the DIEGO method that strategically rectifies key segments of those chains, offers a viable way to overcome this limitation.
@@ IX Solving Generative RPM Problems
Generative RPM problems represent an advanced variant of RPM tasks. In these problems, the solving models are required to generate at least one solution that conforms to the progressive patterns within the given context matrix, rather than selecting from a predefined bounded option pool, thereby more accurately reflecting the model’s comprehension of the reasoning logic. As solution identification remains progressive-pattern-driven, generative RPM problems inherit the multi-solution characteristic from conventional RPM problems. Consequently, models are expected to produce multiple solutions that are structurally diverse, which poses significant challenges for most traditional neural networks.
IX-A WORLD Guides DIO on Generative RPM problems
Beyond enabling DIO to solve discriminative RPM tasks, the WORLD method also equips it with the capability to tackle generative RPM problems. This method constructs a complete generation pipeline for DIO: “estimating solution distribution $\rightarrow$ sampling solution distribution $\rightarrow$ decoding samples”.
The WORLD method extends DIO by incorporating an external GMM $\{U_{m}|m\in[1,M]\}$, an additional loss term $\ell_{\text{Weighted}}$, and a decoder $\mathcal{D}(\cdot)$. Once trained in this framework, DIO can not only discriminate among the options for each context matrix but also directly generate the corresponding solutions. Specifically, as shown in Formula (42), the loss term $\ell_{\text{Weighted}}$ prompts DIO to assign a context-dependent weight to each GMM component $U_{m}$, where the weight reflects the likelihood that $U_{m}$ contains the correct options for the given context matrix. Within this loss term, the weight for each $U_{m}$ given the context matrix features $\{z_{ij}|i\in[1,8]\}$ is calculated as:
$S_{mj}=\tilde{\text{DIO}}(\{z_{ij},U_{m}|i\in[1,8]\}).$ (44)
Since these weights $\{S_{mj}|m\in[1,M],j\in[1,N]\}$ indicate the likelihood that the components cover the correct options, we not only gain the chance to sample diverse incorrect options from low-weight components to tighten the lower bound on the mutual information $I(\{x_{i}\}_{i=1}^{8};x_{\alpha})$, but also have the opportunity to combine high-weight components into a solution distribution. To this end, we adopt the Gumbel-Max trick [37] to sample $N$ high-weight components as the solution distribution that DIO estimates for the context matrix $\{x_{i}|i\in[1,8]\}$. Specifically, this trick is applied as follows:
@@ IX-A WORLD Guides DIO on Generative RPM problems
In this case, the sampled indices $\{\hat{m}_{j}|j\in[1,N]\}$ can point out the solution distribution $\{U_{\hat{m}_{j}}|j\in[1,N]\}$. Subsequently, Formula (39) shows that the well-trained decoder $\mathcal{D}(\cdot)$ is capable of transforming samples from this estimated distribution $\{U_{\hat{m}_{j}}|j\in[1,N]\}$ into the generated answer $\hat{x}$:
@@
In essence, WORLD requires DIO to interact with a GMM that fully encodes RPM image features, a demand that far surpasses traditional discriminative RPM tasks, thereby sparking DIO’s capacity to solve the generative RPM problems.
Notably, WORLD is well-aligned with the multi-solution characteristic of generative RPM problems. Formulas (45)–(47) indicate that the Gumbel-Max trick ensures the estimated solution distribution $\{U_{\hat{m}_{j}}|j\in[1,N]\}$ exhibits diversity. Formula (48) further reveals that the solutions decoded from samples $\{U_{\hat{m}_{j}}+\epsilon_{ij}|j\in[1,N]\}$ of the estimated distribution have the potential to incorporate the sampling diversity $\epsilon_{ij}$. By iterating the “estimating $\rightarrow$ sampling $\rightarrow$ decoding” pipeline, which can generate diverse solutions, WORLD systematically explores multiple potential solutions and fully embraces the multi-solution logic of generative RPM tasks.
IX-B Evaluating the Generated Solutions
Verifying the solution accuracy of generative RPM problems has long been challenging. Generative RPM problems possess the multi-solution characteristic, which means that the generative solving models cannot be evaluated simply by comparing each generated solution with the provided correct option. This also makes it challenging for RPM experts to devise a perfect verification algorithm. Most prior work on generative RPM problems has used self-verification, i.e., using their own discriminative models to evaluate their own generative models. However, this paper argues that this self-verification paradigm is inadequate. In generative RPM problems, the multi-solution characteristic stems from the progressive-pattern-driven solution evaluation approach. This implies that the algorithms or models used to validate the accuracy of solutions for such problems should also be progressive-pattern-driven rather than mutual-information-driven. The DIEGO approach meets this expectation, as the DIO model trained solely with the loss term $\ell_{\text{DIEGO}}$ will be a progressive-pattern-driven solving model.
More importantly, WORLD is a strong mutual-information-driven inductive model. Using the progressive-pattern-driven model DIEGO to verify WORLD’s ability to solve generative RPM problems not only avoids the imprecision of symmetric verification but also examines the rationality and potential of mutual information as a reasoning cue.
@@ X Experiment
All experiments conducted in this paper were programmed in Python based on the PyTorch[38] framework. This paper proposes four distinct configurations for RPM problems:
X-1 Solo DIO
which solves RPM problems through a structured causal chain modeling approach without additional optimization strategies. In the DIO model, the hyperparameter $N$ is set to 16, indicating that the ViT within its image feature extraction module employs a $4\times 4$ patch division strategy;
X-2 Brando-enhanced DIO
which maps out constructive hypothetical incorrect options to tighten the variational lower bound of mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$. In the Brando framework, the hyperparameter $\beta$ is set to 24, which means the Brando network provides 88 ($\beta-16=8$) constructive hypothetical incorrect options to the DIO in each iteration;
X-3 WORLD-augmented DIO
utilizing a Gaussian Mixture Model for feature distribution modeling to provide an infinite set of weighted incorrect options for tightening the lower bound of $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$. In the WORLD framework, the hyperparameter $M$ is set to 256, equipping the GMM employed for DIO with 256 components. This configuration enables the generation of $M^{N}=256^{16}$ distinct solution distributions, theoretically allowing for the sampling of an infinite diversity of incorrect options;
X-4 DIEGO-guided DIO
incorporating metadata supervision to directly rectify semantic biases in the causal chain of abstract attribute patterns. In the DIEGO framework, the parameter $F$ is determined by the number of metadata modality in the targeted RPM problem.
@@ X-A Experiment on RAVEN and I-RAVEN Problems
These results indicate that all four model configurations proposed in this study outperform existing baselines on both the RAVEN and I-RAVEN datasets. Notably, the DIO+DIEGO configuration achieves a peak accuracy of 99.9%, delivering significant performance gains over prior state-of-the-art models (Triple-CFN[33] and RS-TRAN[31]). Particularly in complex reasoning tasks such as $\times3$ Grid and O-IG, the DIO series models demonstrate distinct advantages, validating the generalizability and robustness of their architectural design for abstract visual reasoning challenges.